OpenStudy (dan815):

what is the expected value for throwing m dice, if they have n sides? All m dice are thrown at once, m and n are constants and positive integers THE DICE ARE FAIR!

OpenStudy (freethinker):

use the QH @dan815

OpenStudy (anonymous):

@TuringTest solve his weird question thingy

OpenStudy (anonymous):

lel

OpenStudy (freethinker):

how many times will you be throwing it?

OpenStudy (dan815):

m dice are thrown at once

OpenStudy (ikram002p):

Ok is n a constant value ??

OpenStudy (anonymous):

Why do you and ganeshie always ask pointless no answer questions? ._.

OpenStudy (freethinker):

it is not pointless!

OpenStudy (dan815):

bahahaha I do have a point to this question i am working on something

OpenStudy (ikram002p):

And one more question is n give one side per throwing ??

OpenStudy (dan815):

n sides, so the numbers on each side is 1,2,3... to n

OpenStudy (anonymous):

Is it unbiased?

OpenStudy (freethinker):

@khanacademy

OpenStudy (anonymous):

His username is actually @saifoo.khan

OpenStudy (ikram002p):

And hey lol dan expected value ? I don't get what is that , what else expect numbers ?

OpenStudy (dan815):

like for exampel 2 dice with 6 sides have an expected value of 7

OpenStudy (dan815):

(1*2+2*3+3*4+4*5+5*6+6*7+5*8+4*9+3*10+2*11+1*12 )/36=7

OpenStudy (anonymous):

The average

OpenStudy (dan815):

would it always just be the median i guess?

OpenStudy (anonymous):

mean=median when its a normal distribution

OpenStudy (dan815):

if there are m dice with n sides m*n/2 or if odd floor( m*n/2) +1

OpenStudy (dan815):

LEts see then for 2 dice and 6 sides floor(6*2/2) +1 =7 for 2 dice and 4 sides 4*2/2 +1 =5 1 2 3 4 1 5 2 5 3 5 45

OpenStudy (dan815):

3 dice and 4 sides 4*3/2 +1 = 7 sums 1 1 1 -- 3 1 1 2 ---4 1 2 1 2 1 1 1 1 3 -----5 1 3 1 3 1 1 1 2 2 2 1 2 2 2 1 1 1 4 ------6 1 4 1 4 1 1 . 1 2 3 = 3! ways of this ... 2 2 2 --------7 sum 1 2 4 = 3! ways of this 1 3 3 = 3!/2 2 2 1 = 3!/2 -----------8 ssum 1 3 4 ---- 3! ways 2 3 3 ----3 ways so symmetry along 7, it is probably the expected value

OpenStudy (anonymous):

I think the main questions to answer here is, how many distinct possible outcomes do we have in the general scenario? By definition of expected value, we have \[\large \mathbb{E}(x)=\sum_{x\in X}xp(x)\]It's clear that \(p(x)=\dfrac{1}{n^m}\) for any given outcome. Now it's a matter of determining the cardinality of the sample space \(X\).

OpenStudy (dan815):

Right someone was telling me that we would just consider the median because, we know that after an infinite number of throws we must have a binomial distribution

OpenStudy (dan815):

hey what does cardinality mean

OpenStudy (anonymous):

Size of the set, or number of elements. Basically the total number of outcomes.

OpenStudy (anonymous):

It comes down to finding a proper closed form for the upper limit of the sum.

OpenStudy (dan815):

okay but, how come you are saying p(x) = 1/n^m for any outcome should it be like p(x)= a/n^m, a <n^m

OpenStudy (dan815):

for example the number 3, from throwing 2 dice and 6 sides that has a prob of 2/6^2

OpenStudy (anonymous):

I'm interpreting each outcome to be distinct. Let's say \(m=3\) and \(n=2\). We have \(9\) total outcomes, including repreats: \[\begin{matrix}111&\color{red}{112}&122&222\\ &\color{red}{121}&212\\ &\color{red}{211}&221\end{matrix}\]We have \(2^3=8\) total outcomes, each with \(\dfrac{1}{8}\) probability of occurring, but this doesn't take into account that all the red ones are the same.

OpenStudy (dan815):

ahh i gotcha!

OpenStudy (anonymous):

Whoops, meant \(8\) not \(9\), but yeah

OpenStudy (dan815):

:)